# Ansible Lightspeed: A Code Generation Service for IT Automation

### v1

@@ 3.2. Do users continue to use the service after trying it?
    Users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to accept the suggestion for a thorough examination only to realize it was inaccurate, which made them delete the suggestion.
@@
    Users experimenting with the prompt or examining the deterministic nature of the model, resulting in multiple successive requests after discarding the received suggestion, to see if the answers vary.
The module holds significant importance in an Ansible task. Our analysis reveals that when less than 50% of edits were made, there were 296 instances where users edited the module after accepting the suggestion. This pattern underscores situations where the model’s suggestion did not align with users’ specific module requirements.
So if we consider as a form of rejection the cases where the module is changed or more than 50% of the suggestion is edited, or where the suggestions were deleted after they were first accepted, the actual acceptance rate reduces to 36.05% (down from the initial 48.62%).
We compared our acceptance rate with the acceptance rates of Google Machine Learning enabled code completion system (Maxim Tabachnyk, 2022) (MLECC) which is an internal tool at Google, Github CoPilot (Ziegler et al., 2022) and CodeCompose (Murali et al., 2023), which is an internal tool at Meta. These were the only other code completion tools we found to have used acceptance rate as a metric. Tab. 4 presents the comparison. Google MLECC (Maxim Tabachnyk, 2022) provides acceptance rates for single-line and multi-line code completion, for about 10k internal users, for 8 programming languages over 3 months. They provide user counts for both single-line and multi-line code completion, but don’t breakdown acceptance rate based on languages.
Ziegler et al. (2022) show Co-pilot acceptance rate for survey participants and for all users, for multiple programming languages. Acceptance rate for survey participants are always higher compared to all users, which is why we only show survey participant acceptance rates in Tab. 4. A total of 17,420 participated in their survey. They don’t share total users for each language.
Murali et al. (2023) show CodeCompose acceptance rates and user counts overall and split across multiple languages. We picked top 3 languages with highest acceptance rates. Ansible Lightspeed generates multi-line Ansible code suggestions.
Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and after discarding suggestions which had more than 50% edits.
@@
single-line $C$ 25% 10k+
multi-line $C$ 34% 5k+
@@
Initial accept (Ansible) 48.6% 3,474
Edit¡50% (Ansible) 36.1% 3,474
Finally, to better understand the users behavior, we also study which part of the code the users are editing when they do only small modification. Our findings are as follows:
  1. (1)
    1,503 times out of the 5,247 suggestions with minor edits, users made changes in both the “KEY” and the ”VALUE” parts of the suggested options of the module, added a pair option and value, or removed one. This kind of edit either changes the behavior of the task while staying within the context of the suggested module.
  2. (2)
    In 3,730 instances of minor edits, users opted to change only the “VALUE” component of the received suggestions. This involves personalized modifications to variable names, numeric values, and strings, aligning the suggestions precisely with their unique requirements.
  3. (3)
    We found only 14 such cases in the minor edits category, where the users edited only the ”KEY” component of the received task suggestion. Typically, these occurrences occur when a module has an option with multiple aliases, and the user decides to update an option with its alternative name.
3.4. How do users rate their experience using the explicit feedback mechanism?
Figure 7. Ansible Lightspeed Feedback User Interface.
In VS Code Editor, we have created a dedicated user interface (UI) in the sidebar for Ansible Lightspeed feedback, as shown in Figure 7. Within this UI, users can provide ratings on a scale of 1 to 5, along with written feedback on their experience using the service. We have received a total of 605 user ratings. The user sentiment analysis is performed on this data.
Star rating
@@ 3.4. How do users rate their experience using the explicit feedback mechanism?
The rating in term of stars is presented Figure 8. The data reveals a high level of user satisfaction with the service and recommendations from the Ansible Lightspeed service. Notably, 56.8% of users express considerable satisfaction with the task recommendations, as indicated by their assignment of 4 and 5-star ratings.
However, the sentiment landscape becomes more nuanced when we examine lower rating categories. While the 3-star average rating of 15.9% indicates a moderate level of satisfaction, it’s important to note that users within this category find some value in the recommendations, representing a positive aspect. They recognize areas for improvement but are only somewhat dissatisfied. Our analysis also reveals a 27.3% average rating distribution encompassing 2-star and 1-star ratings. These ratings signify dissatisfaction with the suggestions. Although negative, their feedback serves as constructive input for refining and enhancing the model.
User feedback
In addition to the star rating, we did a thorough analysis of the textual users’ feedback for the Ansible Lightspeed service. We made a manual topic extraction and classification of the of the negative feedback (1 or 2 stars) and positive feedback (3 or 4 stars). The results are presented in Figures 9 and 10.
To provide an impression of the comments, we present a selection of feedback obtained from Ansible Lightspeed users. These quotes have been rephrased for clarity and anonymity, removing any references to internal details.
@@
The majority of the negative feedback, 66.49%, is coming from users that do not manage to get Ansible Lightspeed to work. These problems stem from various sources, including installation difficulties, authentication troubles, network problems, and bugs.
The second largest source of complaints, accounting for 15.71%, is due to poor suggestions provided by the model. This is often caused by the use of uncommon modules, as illustrated by the following example of feedback:
@@
Users are also disappointed with the suggestions provided by the model because they are requesting tasks that fall outside of its current scope. For instance, they may ask for multi-task or playbook generation, but the model is only capable of generating single tasks:
@@
We’re working to address user concerns by continuously improving the service’s documentation and robustness. We’re also expanding the model’s capabilities to handle more use cases and providing customization options so users can choose their preferred collections or provide their own.
@@
On the positive side, 19.72% of customers praise the service’s user-friendliness, highlighting its simplicity in setup and use. For instance, some users have mentioned the login and on-boarding process as being easy and fast. Additionally, a sizable 33.68% of favorable feedback emphasizes the service’s exceptional suggestion accuracy, which practically translates to recommendations for exact code production that align with users’ automation goals. Users have commented that the model not only recommends the correct module but also effectively utilizes module options and sub-options. Furthermore, 42.70% of users consider improved productivity as Ansible Lightspeed’s key benefit. This increase in productivity is supported by observable results like time savings, accelerated documentation activities, and improved process effectiveness.
To illustrate the positive sentiments from the users, we present a selection their feedback.
After the release of the Ansible Lightspeed service, multiple demo sessions were conducted for users. This is what one of the users has to say after using the service for the first time:
“The session was fantastic and was incredibly straightforward to follow. Although I’ve been coding for over two decades, Ansible is entirely new territory for me. Despite my lack of familiarity with Ansible, the AI made it remarkably user-friendly. Thanks to the service, I could create a playbook without any prior knowledge of Ansible, and I mean absolutely none.”
In this case, the user was new to Ansible, but they had a delightful experience writing their first playbook with the help of Ansible Lightspeed. Several remarks that also came from seasoned playbook developers indicate the usefulness of using Ansible Lightspeed in their daily workflow because of its efficiency.
@@
We also found out that upon using the service, users experimented with the prompts and found it to return the correct suggestions most of the time. This is what one of the users has to say about this:
@@
In addition to these remarks, we also observed interesting side effects, such as increased best practices in playbook writing when developers use Ansible Lightspeed. As Ansible Lightspeed runs Ansible Risk Insights (et al, 2020) and Ansible Lint as post-processing tasks, it ensures that the task suggestions follow the best practices. This is evident in the anecdote mentioned below:
“Ansible Lightspeed is fantastic because I don’t need to constantly remember recommended practices, such as using fully qualified collection names (fqcn) every time. The suggestions it provides align with these best practices. The suggestions help me write cleaner loops and maintain consistent white spaces, making my work much smoother.”
This underscores the significance of the user experience for AI-assisted code generation with Ansible Lightspeed. While the service is seen as valuable for playbook developers in IT automation, some feedback compared it with the sophistication of GPT-4. We are actively working to enhance Ansible Lightspeed for an improved experience for all Ansible playbook developers.
@@ 4. Related Work
A lot of work has also been done on developing coding assistants for IDEs based on LLMs. These would include tools that are internal to a company and are only accessible by internal users and those that are openly available. Murali et al. (2023) analyze CodeCompose, an internal code completion tool at Meta which is based on InCoder LLM (Fried et al., 2022), and show that the acceptance rate is about 22% across 9 programming languages for approximately 16k users. Similarly Maxim Tabachnyk (2022) analyze a code completion tool at Google and measure an acceptance rate of 25-34% over 10k+ Google-internal developers. Svyatkovskiy et al. (2020) introduce and evaluate IntelliCode Compose on multiple programming languages, but they use evaluation based on edit-distance compared to the ground truth and do not mention any user acceptance metrics or the total number of users. None of these works mention user sentiment or user retention over an extended period of time.
The openly available coding assistants would be GitHub Copilot (GitHub, 2021), Tab9 (Tab9, 2018), Replit (Replit, 2016), Amazon CodeWhisperer (Amazon, 2023) and Ansible Lightspeed (RedHat, 2023), among others. Among these, GitHub Copilot has been widely studied since it’s release. Nguyen and Nadi (2022) test Copilot on 33 LeetCode questions in four programming languages. Vaithilingam et al. (2022) perform a more user centred evaluation with 24 users to see how programmers use and perceive Copilot. Ziegler et al. (2022) perform an in-depth study of user acceptance, similar to what we do, but for multiple programming languages and show that the acceptance rate of Copilot suggestions for different user categories, for different programming languages, is approximately 20%-30%. Peng et al. (2023) study the impact of Co-pilot on the speed of programmers and find that AI pair programmers are 55.8% faster in implementing an HTTP server in JavaScript. However they do not provide information about acceptance rate of AI pair programmers. Yetiştiren et al. (2023) perform a comparative study of GitHub Copilot, Amazon CodeWhisperer and ChatGPT (OpenAI, 2022) (sibling model of InstructGPT å(Ouyang et al., 2022)) in terms of code quality metrics like code correctness, code security, code reliability and code maintainability but do not consider acceptance of code by real users. A lot of studies on coding assistants do not measure their performance in terms of acceptance rate. Among those that do, we found that Ansible Lightspeed has the highest initial acceptance rate of 48.6%.
We did not find any user retention figures for any of the existing code completion systems or for the VS Code plugins in the VS Code marketplace. In the absence of more relevant baselines, we compare Lightspeed retention rate with that of Android and iOS apps.
@@ 5. Limitations
We do not perform acceptance rate and edit analysis on all Ansible Lightspeed users. Our Telemetry data includes events from users that may not be Ansible programmers and simply want to try out the system; event data from these users would bias the acceptance rate and edit analysis. To mitigate this issue, we created a cohort of returning users, who used Ansible Lightspeed on two or more days.
Ansible Lightspeed’s acceptance rate is compared with models which are much more general, can be used for multiple programming languages and are likely bigger in terms of number of parameters and the amount of data used for training. These general models also tend to have a higher number of users overall and their performance on Ansible-YAML has not been studied or reported. One of the main purposes of our study was to explore the performance of a domain, language specific model vis-a-vis a more general model, with real users. The total number of users, 3,474, is sizable enough for us to be confident about our analysis.
Additionally, our evaluation of Ansible Lightspeed’s user retention lacked official data for comparisons with other code generation services. To address this gap, we compared its user retention with Android and iOS mobile apps, providing insights into its sustained user engagement. While the comparison methodology may not perfectly align with the domain, it highlights Ansible Lightspeed’s ability to retain users over time.
Despite these considerations, the analysis and insights presented here offer a substantial understanding of Ansible Lightspeed’s effectiveness, user engagement, and potential areas for improvement. Based on this work, we hope to continue investigations into domain-specific code generation tools and their impact on IT automation.
@@ 6. Conclusion and Future work
Pujar et al. (2023) show that a relatively small model fine-tuned on a specific Ansible task generation use case can outperform a much larger model on a benchmark dataset. In this paper we show that the suggestions made by such a specifically fine-tuned model can have relatively higher acceptance rate among a large set of users. We also show that the relatively higher acceptance rate results in better user retention compared to the average for Android and iOS apps and generally positive user feedback.
Currently we support only Ansible task generation, however we are working to expand the capabilities of the model to be able to write entire Ansible playbook based on user prompt. We also plan to develop capabilities for Ansible code explanation, debugging and as well as customization for specific users.

### v2

@@ 3.4. User Acceptance Rate
    The suggestion is initially accepted by the user.
@@
    After acceptance, the suggestion is not deleted.
  3. (3)
    Suggestion has minimal edits or less than 50% of the suggestion is edited.
  4. (4)
    Critical components, like the module name in case of Ansible, should not be edited.
As per this definition, the Edit¡50% acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%. We feel Edit¡50% acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
Table. 2 shows Ansible Lightspeed acceptance rate compared to other IDE based code completion tools. We compared our acceptance rate with the initial acceptance rates of Google Machine Learning enabled code completion system (Maxim Tabachnyk, 2022) (MLECC) which is an internal tool at Google, Github CoPilot (Ziegler et al., 2022) and CodeCompose (Murali et al., 2023), which is an internal tool at Meta. These were the only other code completion tools we found to have used acceptance rate as a metric. Google MLECC (Maxim Tabachnyk, 2022) provides acceptance rates for single-line and multi-line code completion, but don’t breakdown acceptance rate based on languages. Ziegler et al. (2022) for GitHub Co-pilot and Murali et al. (2023) for CodeCompose share acceptance rate for individual languages. We show the top three languages with best acceptance rate. The acceptance rate for these three baselines is what what we call the initial acceptance rate. It merely tells us whether the user accepted the suggestion offered without telling us how the user actually used the suggestion. It does not give any insight into whether the user accepted the suggestion as is, or edited the suggestion or altogether deleted it.
Ansible Lightspeed generates multi-line Ansible code suggestions. Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Edit¡50% acceptance rate.
@@
single-line CC 25% 10k+
multi-line CC 34% 5k+
@@
Initial accept (Ansible) 65.9% 3,910
Edit¡50% (Ansible) 49.1% 3,910
Q: At what rate do users accept suggestions? Initial acceptance rate is misleading since users can delete or heavily modify suggestions. 65.9% of suggestions are initially accepted. We define new acceptance rate, Edit¡50% Acceptance rate. 49.1% of model suggestions were accepted as per Edit¡50% Acceptance rate. Ansible Lightspeed has relatively higher acceptance for Ansible, compared to more general tools for various programming languages.
3.5. Rejected Suggestion Analysis
There are a total of 1,710 initially accepted suggestions which were explicitly edited by the users at the module level. We consider such suggestions to be rejected for Edit¡50% Acceptance rate. However, the users must have found something useful in the suggestion for them to accept it and then spend sometime editing it. We analyse these edits in order to better understand how the model suggestions differ from user expectations.
3.5.1. User preference
In order to remove ambiguity in module names, Ansible Lint encourages users to use Fully Qualified Collection Name, or FQCN (Sbarnea, 2023), when declaring modules. In keeping with this best practice, Ansible Lightspeed always generates module names in FQCN format. On the other hand it is quite possible that some users prefer the shorter version of the module name. For example, the best practice FQCN reference would be ansible.builtin.debug while some users may prefer to refer to the same module as debug.
We find that in 346, or 20.2% of, instances the users edited the model suggestion and changed the FQCN module name to it’s shortened format. The body of the generated Ansible task often had little to no edits, indicating that the FQCN module name, which is correct, is not what they prefer.
3.5.2. YAML reorganization
As per our observation, in the process of writing an Ansible playbook users often reorganize the YAML structure. This reorganization involves adding new valid Ansible keys like block, tag, register, loop and become. The model generated suggestion is generally part of the reorganized playbook, either in the edited or original format, and is likely to have inspired the reorganization.
Such edits indicate that the generated suggestion is not necessarily incorrect, but is definitely incomplete and can be better organized. There are at least 211 (12.3%) such instances of reorganization.
3.5.3. Functionally similar modules
Module confusion usually occurs when there are multiple functionally similar modules which can be used to implement user’s intent and the user is at least somewhat familiar with one of them.
One example of this is command and shell modules, which are often interchangeable. In addition, functionalities of many other modules can be implemented by command or shell modules. It is possible that users are not aware of other modules, or they may prefer another module than the one that lightspeed generated. Generated shell commands are also likely to be incorrect, which could inspire more user edits. There are 352 edits (20.6%) which are related to the command/shell module. Apart from command or shell module related edits, we found 336 (19.6%) instances of edits related to modules which have similar functionalities and can be used interchangeably.
There are another 586 (28%) instances where the user edited the generated module. Understanding these edits requires further analysis.
Q: Why do users edit module names in accepted suggestions? Suggestion is edited to remove FQCN based on user preference. YAML is reorganized to add more details which the model suggestion is missing. Certain functionalities can be implemented in multiple ways. Users change the suggestion based on their preferred implementation.
4. User Feedback
A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor, as shown in Figure 4.
4.1. Star rating
@@
[…]
A breakdown of ratings in terms of stars is presented in Figure 9. Notably, a total of 57% of users express considerable satisfaction with the service, as indicated by their assignment of 4 and 5-star ratings. The 3-star rating by 15.8% of users indicates a moderate level of satisfaction. In this case users do find some value in using Ansible Lightspeed, but expect more improvement. About 27.2% users appeared dissatisfied and gave a rating of 2 and 1-star.
4.2. User Comments
Apart from star-rating, users also submit feedback in the form of textual comments. By analyzing these comments we hope to find areas for improvement and also know what users like about the service. We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars) into various reasons. We also share some examples of real user comments, both positive and negative. These quotes have been rephrased for clarity and anonymity, removing any references to user details.
@@ 4.2. User Comments
[…]
The breakdown of negative comments based on manual review is presented in Figure 10. The majority of the negative feedback, 66.49%, is coming from users that do not manage to get Ansible Lightspeed to work. These problems stem from various sources, including installation difficulties, authentication troubles, network problems, and bugs.
The second largest source of complaints, accounting for 15.71%, is due to poor suggestions provided by the model. This could sometimes be caused by the use of uncommon modules, as illustrated by the following example of feedback:
@@
Some users are disappointed with the suggestions provided by the model because their requests fall outside of Lightspeed’s current scope. For instance, they may ask for multi-task or playbook generation, but the model is currently only trained for generating single tasks:
@@
[…]
Breakdown of reasons for positive experience is presented in Figure 11. Most users, about 42.7%, consider improved productivity as Ansible Lightspeed’s key benefit. Multiple users mentioned that they noticeably saved time by using Lightspeed.
@@
33.7% of users found suggestion accuracy to be the most impressive part of Ansible Lightspeed. Users have commented that the model not only recommends the correct module but also the associated module options and values.
@@
About 19.7% of customers praise the service’s user-friendliness, highlighting its simplicity in setup and use. For instance, some users have mentioned the login and on-boarding process as being easy and fast. Remaining 3.9% of the users said best part was the general experience of using the tool.
@@ 5. Related Work
A lot of work has also been done on developing coding assistants for IDEs based on LLMs. These would include tools that are internal to a company and are only accessible by internal users and those that are openly available. Murali et al. (2023) analyze CodeCompose, an internal code completion tool at Meta which is based on InCoder LLM (Fried et al., 2022), and show that the acceptance rate is about 22% across 9 programming languages for approximately 16k users. Similarly, Maxim Tabachnyk (2022) analyze a code completion tool at Google and measure an acceptance rate of 25-34% over 10k+ Google-internal developers. Svyatkovskiy et al. (2020) introduce and evaluate IntelliCode Compose on multiple programming languages, but they use evaluation based on edit-distance compared to the ground truth and do not mention any user acceptance metrics or the total number of users. None of these works mention user sentiment or user retention over an extended period of time.
The openly available coding assistants would be GitHub Copilot (GitHub, 2021), Tab9 (Tab9, 2018), Replit (Replit, 2016) and Amazon CodeWhisperer (Amazon, 2023) among others. Among these, GitHub Copilot has been widely studied since its release. Nguyen and Nadi (2022) test Copilot on 33 LeetCode questions in four programming languages. Vaithilingam et al. (2022) perform a more user-centred evaluation with 24 users to see how programmers use and perceive Copilot. Ziegler et al. (2022) perform an in-depth study of user acceptance, similar to what we do, but for multiple programming languages and show that the acceptance rate of Copilot suggestions for different user categories, for different programming languages, is approximately 20%-30%. Peng et al. (2023) study the impact of Co-pilot on the speed of programmers and find that AI pair programmers are 55.8% faster in implementing an HTTP server in JavaScript. However, they do not provide information about the acceptance rate of AI pair programmers. Yetiştiren et al. (2023) perform a comparative study of GitHub Copilot, Amazon CodeWhisperer and ChatGPT (OpenAI, 2022) (sibling model of InstructGPT (Ouyang et al., 2022)) in terms of code quality metrics like code correctness, code security, code reliability and code maintainability but do not consider acceptance of code by real users.
We did not find any user retention figures for any of the existing code completion systems or for the VS Code plugins in the VS Code marketplace.
@@ 6. Limitations
We do not perform acceptance rate and edit analysis on all Ansible Lightspeed users. Our Telemetry data includes events from users who may not be Ansible programmers and simply want to try out the system with toy examples, which could add noise to our analysis. To mitigate this issue, we created a cohort of returning users who used Ansible Lightspeed on two or more days.
Ansible Lightspeed’s acceptance rate is compared with models which are much more general, can be used for multiple programming languages and are likely bigger in terms of number of parameters and the amount of data used for training. These general models also tend to have a higher number of users overall and their performance on Ansible-YAML has not been studied or reported. One of the main purposes of our study was to explore the performance of a domain, language-specific model vis-a-vis a more general model with real users. The total number of users, 3,910, is sizable enough for us to be confident about our analysis.
Additionally, our evaluation of Ansible Lightspeed’s user retention lacked official data for comparisons with other code generation services. Lacking user retention baseline for code completion tools, we can’t say with confidence that users find Ansible Lightspeed useful enough to keep using it, especially for work. This is why we rely on temporal trends, acceptance rates, and explicit feedback to make such assessments.
Despite these considerations, we hope the analysis and insights presented here offer a substantial understanding of Ansible Lightspeed’s effectiveness, user engagement, and potential areas for improvement. Based on this work, we hope to continue investigations into domain-specific code generation tools and their impact on IT automation.
@@ 7. Conclusion and Future work
Pujar et al. (2023) show that a relatively small model fine-tuned on a specific Ansible task generation use case can outperform a much larger model on a benchmark dataset. In this paper, by analyzing user interaction data of 3,910 users, we show that the suggestions made by such a specifically fine-tuned model can have a relatively higher acceptance rate among a large set of users.
Analysis of feedback from 605 users and of 1,710 user edits to Lightspeed generated Ansible tasks led to interesting findings which we plan to use to improve Ansible Lightspeed. Currently, we use greedy decoding to generate one suggestion for every prompt. To account for user preferences and the fact that there can be multiple modules that can be used to implement similar functionalities, we hope to generate multiple suggestions. Also, currently we support only Ansible task generation, but we are working to expand the capabilities of the model to be able to generate longer Ansible sequences. With this, we hope to reduce the need for the user to reorganize the code and add missing details. We also plan to develop capabilities for Ansible code explanation, debugging, and as well as customization for specific users.
We hope that our proposed user edit-based acceptance criteria are used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful. We also hope that by sharing Ansible Lightspeed’s user retention figures, others get encouraged to share similar figures for their code completion tools. And finally, by sharing the code for Ansible lightspeed service [^7] https://github.com/ansible/ansible-ai-connect-service and our analysis framework [^8] https://github.com/ansible-community/lightspeed-analysis-framework we hope to help others trying to develop similar tools.